# The numeraire e-variable

### 29 Feb 2024 v1

@@ Appendix A A generalized Lebesgue decomposition
The following generalized Lebesgue decomposition into a ‘regular’ and ‘singular’ part is used to define the extended numeraire in Section 2. As always, we work on a given measurable space $(\Omega,{\mathcal{F}})$. Given a nonempty subset ${\mathcal{P}}\subset M_{+}$, an event $N\in{\mathcal{F}}$ is called ${\mathcal{P}}$-negligible if ${\mathbb{P}}(N)=0$ for all ${\mathbb{P}}\in{\mathcal{P}}$.
@@
For any ${\mathbb{Q}}\in M_{+}$ and nonempty ${\mathcal{P}}\subset M_{+}$ there is a unique decomposition ${\mathbb{Q}}={\mathbb{Q}}^{r}+{\mathbb{Q}}^{s}$ with ${\mathbb{Q}}^{r},{\mathbb{Q}}^{s}\in M_{+}$ such that ${\mathbb{Q}}^{r}\ll{\mathcal{P}}$ and ${\mathbb{Q}}^{r},{\mathbb{Q}}^{s}$ are singular, meaning that there exists a ${\mathcal{P}}$-negligible event $N\in{\mathcal{F}}$ such that ${\mathbb{Q}}^{s}(N^{c})=0$. The event $N$ is essentially unique, i.e., unique up to ${\mathbb{Q}}$- and ${\mathbb{P}}$-nullsets for all ${\mathbb{P}}\in{\mathcal{P}}$. In particular, any other ${\mathcal{P}}$-negligible event $N^{\prime}$ satisfies ${\mathbb{Q}}(N^{\prime}\setminus N)=0$.
@@
Define $\alpha=\sup_{N\in{\mathcal{N}}}{\mathbb{Q}}(N)$ where ${\mathcal{N}}=\{N\in{\mathcal{F}}\colon{\mathbb{P}}(N)=0\text{ for all }{\mathbb{P}}\in{\mathcal{P}}\}$ is the family of all ${\mathcal{P}}$-negligible sets. Note that ${\mathcal{N}}$ is closed under countable unions. Pick a maximizing sequence $N_{n}$ and set $N=\bigcup_{n}N_{n}$. Then $\alpha\geq{\mathbb{Q}}(N)\geq{\mathbb{Q}}(N_{n})\to\alpha$, so ${\mathbb{Q}}(N)=\alpha$. Now define ${\mathbb{Q}}^{s}(A)={\mathbb{Q}}(A\cap N)$ and ${\mathbb{Q}}^{r}(A)={\mathbb{Q}}(A\cap N^{c})$. To see that ${\mathbb{Q}}^{r}\ll{\mathcal{P}}$, suppose this is not the case. Then there is some $N^{\prime}\in{\mathcal{N}}$ with ${\mathbb{Q}}^{r}(N^{\prime})>0$, and hence ${\mathbb{Q}}(N\cup N^{\prime})\geq{\mathbb{Q}}^{s}(N)+{\mathbb{Q}}^{r}(N^{\prime})>\alpha$, a contradiction. To show uniqueness, suppose we have two tuples ${\mathbb{Q}}^{r}_{i},{\mathbb{Q}}^{s}_{i},N_{i}$ for $i=1,2$ with the stated properties. Set $N=N_{1}\cup N_{2}\in{\mathcal{N}}$ and note that ${\mathbb{Q}}^{s}_{i}(N^{c})\leq{\mathbb{Q}}^{s}_{i}(N_{i}^{c})=0$. Then for any $A$, we have ${\mathbb{Q}}(A\cap N)={\mathbb{Q}}^{s}_{i}(A\cap N)={\mathbb{Q}}^{s}_{i}(A)$. Thus ${\mathbb{Q}}^{s}_{1}={\mathbb{Q}}^{s}_{2}$, and then also ${\mathbb{Q}}^{r}_{1}={\mathbb{Q}}^{r}_{2}$. Furthermore, the symmetric difference $N_{1}\Delta N_{2}$ is a ${\mathbb{Q}}$-nullset because ${\mathbb{Q}}(N_{1}\cap N_{2}^{c})={\mathbb{Q}}_{2}^{r}(N_{1})=0$ and similarly ${\mathbb{Q}}(N_{2}\cap N_{1}^{c})=0$. Since also $N_{1}\Delta N_{2}$ belongs to ${\mathcal{N}}$, the essential uniqueness statement follows. Finally, if $N^{\prime}\in{\mathcal{N}}$ then ${\mathbb{Q}}(N^{\prime}\setminus N)={\mathbb{Q}}^{r}(N^{\prime}\setminus N)+{\mathbb{Q}}^{s}(N^{\prime}\setminus N)\leq{\mathbb{Q}}^{r}(N^{\prime})+{\mathbb{Q}}^{s}(N^{c})=0$. ∎

### 22 Mar 2024 v2

@@ Appendix A A generalized Lebesgue decomposition
The following generalized Lebesgue decomposition into a ‘regular’ and ‘singular’ part is used in the proof of Lemma 2.8. As always, we work on a given measurable space $(\Omega,{\mathcal{F}})$. Given a nonempty subset ${\mathcal{P}}\subset M_{+}$, an event $N\in{\mathcal{F}}$ is called ${\mathcal{P}}$-negligible if ${\mathsf{P}}(N)=0$ for all ${\mathsf{P}}\in{\mathcal{P}}$.
@@
For any ${\mathsf{Q}}\in M_{+}$ and nonempty ${\mathcal{P}}\subset M_{+}$ there is a unique decomposition ${\mathsf{Q}}={\mathsf{Q}}^{r}+{\mathsf{Q}}^{s}$ with ${\mathsf{Q}}^{r},{\mathsf{Q}}^{s}\in M_{+}$ such that ${\mathsf{Q}}^{r}\ll{\mathcal{P}}$ and ${\mathsf{Q}}^{r},{\mathsf{Q}}^{s}$ are singular, meaning that there exists a ${\mathcal{P}}$-negligible event $N\in{\mathcal{F}}$ such that ${\mathsf{Q}}^{s}(N^{c})=0$. The event $N$ is essentially unique, i.e., unique up to ${\mathsf{Q}}$- and ${\mathsf{P}}$-nullsets for all ${\mathsf{P}}\in{\mathcal{P}}$. In particular, any other ${\mathcal{P}}$-negligible event $N^{\prime}$ satisfies ${\mathsf{Q}}(N^{\prime}\setminus N)=0$.
@@
Define $\alpha=\sup_{N\in{\mathcal{N}}}{\mathsf{Q}}(N)$ where ${\mathcal{N}}=\{N\in{\mathcal{F}}\colon{\mathsf{P}}(N)=0\text{ for all }{\mathsf{P}}\in{\mathcal{P}}\}$ is the family of all ${\mathcal{P}}$-negligible sets. Note that ${\mathcal{N}}$ is closed under countable unions. Pick a maximizing sequence $N_{n}$ and set $N=\bigcup_{n}N_{n}$. Then $\alpha\geq{\mathsf{Q}}(N)\geq{\mathsf{Q}}(N_{n})\to\alpha$, so ${\mathsf{Q}}(N)=\alpha$. Now define ${\mathsf{Q}}^{s}(A)={\mathsf{Q}}(A\cap N)$ and ${\mathsf{Q}}^{r}(A)={\mathsf{Q}}(A\cap N^{c})$. To see that ${\mathsf{Q}}^{r}\ll{\mathcal{P}}$, suppose this is not the case. Then there is some $N^{\prime}\in{\mathcal{N}}$ with ${\mathsf{Q}}^{r}(N^{\prime})>0$, and hence ${\mathsf{Q}}(N\cup N^{\prime})\geq{\mathsf{Q}}^{s}(N)+{\mathsf{Q}}^{r}(N^{\prime})>\alpha$, a contradiction. To show uniqueness, suppose we have two tuples ${\mathsf{Q}}^{r}_{i},{\mathsf{Q}}^{s}_{i},N_{i}$ for $i=1,2$ with the stated properties. Set $N=N_{1}\cup N_{2}\in{\mathcal{N}}$ and note that ${\mathsf{Q}}^{s}_{i}(N^{c})\leq{\mathsf{Q}}^{s}_{i}(N_{i}^{c})=0$. Then for any $A$, we have ${\mathsf{Q}}(A\cap N)={\mathsf{Q}}^{s}_{i}(A\cap N)={\mathsf{Q}}^{s}_{i}(A)$. Thus ${\mathsf{Q}}^{s}_{1}={\mathsf{Q}}^{s}_{2}$, and then also ${\mathsf{Q}}^{r}_{1}={\mathsf{Q}}^{r}_{2}$. Furthermore, the symmetric difference $N_{1}\Delta N_{2}$ is a ${\mathsf{Q}}$-nullset because ${\mathsf{Q}}(N_{1}\cap N_{2}^{c})={\mathsf{Q}}_{2}^{r}(N_{1})=0$ and similarly ${\mathsf{Q}}(N_{2}\cap N_{1}^{c})=0$. Since also $N_{1}\Delta N_{2}$ belongs to ${\mathcal{N}}$, the essential uniqueness statement follows. Finally, if $N^{\prime}\in{\mathcal{N}}$ then ${\mathsf{Q}}(N^{\prime}\setminus N)={\mathsf{Q}}^{r}(N^{\prime}\setminus N)+{\mathsf{Q}}^{s}(N^{\prime}\setminus N)\leq{\mathsf{Q}}^{r}(N^{\prime})+{\mathsf{Q}}^{s}(N^{c})=0$. ∎